A collection of nickels, dimes, and quarters consist of 10 coins with a total of $1.10. If the number of dimes is equal to the number of nickels, find the number of each type of coins. Please help solve with steps, thanks. :)
Found 2 solutions by macston, lwsshak3:
Answer by macston(5194)

.
N=number of nickels; D=number of dimes; Q=number of quarters
.
D=N
N+D+Q=10
N+N+Q=10
2N+Q=10
Q=10-2N
.
$0.05N+$0.10D+$0.25Q=$1.10
$0.05N+$0.10N+$0.25Q=$1.10
$0.15N+$0.25Q=$1.10
$0.15N+$0.25(10-2N)=$1.10
$0.15N+$2.50-$0.50N=$1.10
-$0.35N=-$1.40
N=4
ANSWER 1: There are 4 nickels.
.
D=N+4
ANSWER 2: There are 4 dimes.
.
Q=10-2N=10-2(4)=10-8=2
ANSWER 3: There are 2 quarters.
.
CHECK:
$0.05N+$0.10D+$0.25Q=$1.10
$0.05(4)+$0.10(4)+$0.25(2)=$1.10
$0.20+$0.40+$0.50=$1.10
$1.10=$1.10


Answer by lwsshak3(11628)

A collection of nickels, dimes, and quarters consist of 10 coins with a total of $1.10. If the number of dimes is equal to the number of nickels, find the number of each type of coins. Please help solve with steps, thanks. :)
let x=number of dimes and nickels
10-2x=number of quarters
.05x+.10x+.25(10-2x)=1.10
.05x+.10x+2.5-.5x=1.10
.35x=1.40
x=4
10-2x=10-8=2
number of dimes and nickels=4
10-2x=number of quarters=2